A 25 ounce solution is 20 percent alcohol. If 50 ounces of water are alcohol to it, what percent of the new solution is alcohol?'
Answer by richwmiller(17219)

25*.80+50*1.0=75w
20+50=75w
70=75w
14/15=w
0.933=water
5/75=.06667%=alcohol
or

25*.20=75a
5=75a
5/75=.06667%=alcohol